How do i use distributive property to write expression without parentheses and then simplify? on these two problems
-( k -3 + 3m) + 3
and
-.6(3p +1)-.6

Answer by nyc_function(260)

-( k -3 + 3m) + 3 = -k + 3 - 3m


-.6(3p +1)-.6 = -1.8p - 0.6 - 0.6 = -1.8p - 1.2